OpenStudy (brightemmerald):

20x^2-26x+8=0 Which of the below are equal to 0? Using the equation above, check all that apply. x=2/3 x=1/5 x=3/5 x=1/2 x=1/3 x=4/5

OpenStudy (princeharryyy):

try replacing x =1/2 in the equation

OpenStudy (will.h):

you need to find the zeros? if so you will have to use \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (eliesaab):

Solve the quadratic equation to get x=1/2 or x=4/5

OpenStudy (eliesaab):

Use the quadratic formula

OpenStudy (will.h):

@eliesaab share your steps don't give direct answers

OpenStudy (brightemmerald):

We have to turn the equation into a binomial, then try to turn the binimials = to 0.

OpenStudy (will.h):

have you learn about the famous quadratic formula? the one i posted?

OpenStudy (eliesaab):

Or factor it \[ 20 x^2-26 x+8=2 (2 x-1) (5 x-4) \]

OpenStudy (will.h):

okay so group factoring is pretty good way too 20x^2-26x+8=0 1st of all what is the GCF?

OpenStudy (brightemmerald):

4 is the GCF

OpenStudy (will.h):

great simplify what would be the new shape of the expression?

OpenStudy (brightemmerald):

wait... 26 doesnt have a GCF of 6

OpenStudy (brightemmerald):

err 4

OpenStudy (will.h):

good note so what would be the GCF?

OpenStudy (will.h):

that would help indeed @eliesaab you deserve a medal